OpenStudy (taytay02):

In heptagon Q R S T U V W angle Q measures one hundred thirty degrees, angle R measures one hundred forty degrees, angle S and angle T measure one hundred sixty degrees each, angle U measures one hundred degrees, and angle V measures ninety degrees. What is the measure of angle W?

OpenStudy (taytay02):

is it 130

OpenStudy (anonymous):

is there a picture or no??

OpenStudy (taytay02):

no

OpenStudy (anonymous):

@SolomonZelman Can You Help @taytay02

OpenStudy (taytay02):

somebody

OpenStudy (taytay02):

@SolomonZelman

OpenStudy (anonymous):

@tkhunny can you help @taytay02

OpenStudy (taytay02):

wait no its 120 nevermind i figured it out

OpenStudy (tkhunny):

7 angles. What is the total interior angle measure of a heptagon? I always remember the Exterior Angles Relationships. 360/7 = 51.42857 180-51.42857 = 128.57143 7*128.57143 = 900 Also, (7-2)*180 = 900 -- If you happen to remember the Interior Angle Formula. Well, they have to add up to 900º if it's a heptagon.